CS 154 Winter Quarter 1996: Instructor: Rob van Glabbeek

Grader Return Sheet

Be sure to fill this out, and clip this to the top of the pile of homeworks, when you return your set of homeworks after grading.^{1}

Full name (print): Email:

Homework Number: Problems (list numbers):

Coordinating TA (circle):  Nikolaj  Harish

Time and Day you returned your homeworks:

Here is a checklist of things you should keep track of through the grading process:

• Lookup http://theory.stanford.edu/~cs154/grading-plan.html for the timeline

of your grading involvement.

• Use a red pen to make comments (legibly!), assign points etc. Try not to do scratch work on the students’ answer sheet.

• The final score for each problem should be written clearly and circled. Put your initials next to each final score.

• The grading team for your grading atom has to prepare model solutions for that portion. If you are submitting your team’s model solution, attach a photocopier-ready version of the solutions to this sheet.

• Sort the homeworks in alphabetical order by last name when you turn them in.

• The deadline for receiving graded homeworks, model solutions and the electronic scoresheet is Monday @ noon. This is a hard deadline as grades have to be reviewed, solutions photocopied and homeworks returned to class, all by the same day.

• You may return your papers to Nikolaj (Gates 460), Harish (Gates 490), any of their officemates, Rob (Gates 478) or leave them at Rob’s mail folder (Gates 465) at any time prior to the deadline. If you do not hand in the homeworks personally to course staff, send email to cs154@theory.stanford.edu to inform us.

1A copy of this sheet is also on the course web page.